OpenStudy (anonymous):

A bag contains 12 pieces of fruit: 7 apples, 3 oranges, and 2 nectarines. Jonas is given a piece of fruit first and Beth is given one next. What is the probability that Jonas will get an orange and Beth will get an apple if each is handed a piece of fruit at random?

OpenStudy (anonymous):

first compute the probability that jonas gets an orange it is the ratio of the number of oranges to the total number of pieces of fruit one more step after that

OpenStudy (anonymous):

@satellite73 it would be 3/12 then what?

OpenStudy (anonymous):

That is one BIG bag.

OpenStudy (anonymous):

then compute the probability that beth gets an apple "given" that jonas has an orange we do this by thinking there were originally 12 pieces of fruit, now there are only 11 (since jonas got one) jonas got an orange, so there are still 7 apples in the bag

OpenStudy (anonymous):

therefore the probability that beth gets an apple "given" that jonas has an orange is \(\frac{7}{11}\)

OpenStudy (anonymous):

last step is to take these to numbers and multiply them together

OpenStudy (anonymous):

*two numbers

OpenStudy (anonymous):

ok so it would be 7/44 right?? @satellite73

OpenStudy (anonymous):

\[\frac{1}{4}\times \frac{7}{11}\] yes you are right

OpenStudy (anonymous):

thank you @satellite73